# MeGA: Merging Multiple Independently Trained Neural Networks Based on Genetic Algorithm

### v1

@@ 1 Introduction
Training models in parallel rather than sequentially offers efficiency and practicality. Multiple models can learn diverse patterns quickly, and merging their weights can achieve superior performance faster [15, 3, 11]. This approach is particularly advantageous in distributed learning environments, facilitating scalable and efficient training across multiple GPUs or devices [3, 23, 1].
@@ 3.1.3 Initialization
The weights of each individual $\boldsymbol{\theta}_{i}$ are initialized as a linear combination of the weights from two pre-trained models, $\boldsymbol{\theta}_{1}$ and $\boldsymbol{\theta}_{2}$:
$\boldsymbol{\theta}_{i}=\alpha\boldsymbol{\theta}_{1}+(1-\alpha)\boldsymbol{\theta}_{2},\quad\alpha\sim\text{Uniform}(0,1)$
Here, $\alpha$ is a random scalar drawn from a uniform distribution between 0 and 1. This ensures that each individual’s weights are a unique blend of the two parent models. By initializing the population in this manner, we promote diversity among the candidate solutions, helping the genetic algorithm effectively explore the solution space and avoid premature convergence.
The combination of weights is applied element-wise to preserve the detailed characteristics of both parent models. Let $\theta_{1,j,k}$ and $\theta_{2,j,k}$ denote the $(j,k)$-th element of the weight matrices of the two pre-trained models. The $(j,k)$-th element of the combined weight matrix for the $i$-th individual is given by:
@@ 3.1.5 Selection
Next, the fitness $F(\boldsymbol{\theta})$ of each individual $\boldsymbol{\theta}$ in the tournament set $\mathcal{T}$ is evaluated using the previously defined fitness function:
$F(\boldsymbol{\theta})=\text{Accuracy}(f(\mathbf{X}_{\text{val}};\boldsymbol{\theta}),\mathbf{y}_{\text{val}})$
After evaluating the fitness of all individuals in the tournament set, the individual with the highest fitness is selected as a parent:
@@ 3.1.6 Crossover
Crossover is performed on pairs of selected parents to generate offspring. This process combines the genetic material (weights) of two parents to produce a new individual (offspring), thereby promoting genetic diversity. Given two parents $\boldsymbol{\theta}_{a}$ and $\boldsymbol{\theta}_{b}$, an offspring $\boldsymbol{\theta}_{\text{child}}$ is created by taking a weighted combination of the parents’ weights:
$\boldsymbol{\theta}_{\text{child}}=\beta\boldsymbol{\theta}_{a}+(1-\beta)\boldsymbol{\theta}_{b},\quad\beta\sim\text{Uniform}(0,1)$
The crossover is performed element-wise for each weight:
@@ 4 Experimental Results
Figure 2: MeGA Fitness Progression for ResNet-56 and DenseNet 121 models over 20 generations. The left plot shows the best fitness values for the ResNet-56 model, starting from 0.8042 and improving to 0.8224. The right plot illustrates the best fitness values for the DenseNet 121 model, starting from 0.7290 and improving to 0.7427. These plots demonstrate the effectiveness of the genetic algorithm in optimizing the weight combinations of both models.
@@ 4.1 Experimental Settings
We utilized the CIFAR-10 dataset [19]. We trained Convolutional Neural Network (CNN) models: ResNet [13], Xception [2], and DenseNet [16] without pre-trained weights and without data augmentation. These models were trained using a batch size of 256 over 50 epochs with the Adam optimizer [18], using a learning rate of 0.01. Each model followed the same architecture and hyperparameters to ensure consistency and a fair comparison. After training, we applied the genetic algorithm-based neural network merging (MeGA) approach to merge the weights of these models into a single set of weights. The performance of the merged model was then evaluated and compared against the individual models. For all experiments, we used the following MeGA hyperparameters: a population size of 20, 20 generations, 4 parents per generation, and a mutation rate of 0.02.
4.2 Image Classification Results
The results of the image classification experiments on the CIFAR-10 dataset [19] demonstrate the effectiveness of the Genetic Algorithm-based Neural Network (MeGA) weight merging approach, as shown in Table 2.
@@ 4.2 Image Classification Results
Model (# of Params.) Baseline (Random Seed). Baseline Acc. Final Test Acc. (Single Model) (Merged Model) ResNet 56 (0.86M) [13] ResNet 56 Model 1 (56) 0.809 0.822 ResNet 56 Model 2 (57) 0.793 ResNet 110 (1.78M) [13] ResNet 110 Model 1 (56) 0.801 0.816 ResNet 110 Model 2 (57) 0.814 ResNet 152 (3.51M) [13] ResNet 152 Model 1 (56) 0.783 0.819 ResNet 152 Model 2 (57) 0.781 Xception (22.96M) [13] Xception Model 1 (56) 0.716 0.754 Xception Model 2 (57) 0.730 DenseNet 121 (7.04M) [13] DenseNet 121 Model 1 (56) 0.719 0.742 DenseNet 121 Model 2 (57) 0.728 DenseNet 169 (12.65M) [13] DenseNet 169 Model 1 (56) 0.735 0.753 DenseNet 169 Model 2 (57) 0.731
The improvements in fitness values were reflected in the test accuracies of the models. The ResNet models of varying depths (56, 110, and 152 layers) showed improvements in test accuracy after merging the weights using the MeGA approach. For instance, the ResNet 56 model’s test accuracy improved from an average of 0.801 to 0.822. The ResNet 110 model, which has 1.78 million parameters, demonstrated a final test accuracy of 0.824 after merging, indicating that the MeGA approach effectively leveraged the strengths of the individual models.
Similarly, the Xception model, which uses depthwise separable convolutions, saw an increase in test accuracy from an average of 0.723 to 0.754 after the MeGA-based weight merging. This improvement highlights the MeGA approach’s ability to optimize and combine complex model architectures effectively.
The DenseNet models (121 and 169 layers) also benefited from the MeGA approach, with improvements in test accuracy. The DenseNet 121 model’s accuracy increased from an average of 0.724 to 0.742, while the DenseNet 169 model’s accuracy increased from 0.733 to 0.753.
Overall, the MeGA approach consistently outperformed the individual models in terms of test accuracy. The merged model’s performance was significantly better than that of any single model, demonstrating the potential of the MeGA approach in optimizing and combining neural network weights to achieve superior performance. This suggests that the genetic algorithm-based weight merging technique is a viable and efficient method for enhancing the performance of neural networks by integrating multiple pre-trained models.
@@ 4.3 Extended Experiments for Multi-Models
In this section, we illustrate the application of our genetic algorithm-based weight merging method to combine multiple neural network models. This process leverages the collective strengths of various models, thereby enhancing overall performance. The hierarchical merging approach we employed systematically and effectively merges the weights of eight independently trained ResNet56 models into a single robust model. The hierarchical structure of this merging process is depicted in Figure 3.
The hierarchical merging process begins with training eight ResNet56 [13] models on the CIFAR-10 dataset [19]. These models are then paired and merged using a genetic algorithm, resulting in four intermediate models: Merged Network 1 through Merged Network 4. The algorithm optimizes weight combinations through selection, crossover, and mutation. Next, these four intermediate models are further merged into two higher-level models, Merged Network 5 and Merged Network 6, using the same genetic algorithm approach. Finally, the two higher-level models are merged into a single final model, the Final Network, ensuring it combines the strengths of all eight original models.
@@
Model (# of Params). Baseline (Random Seed). Baseline Test Acc. Final Test Acc. (Single Model) (Merged Model) ResNet 56 (0.86M) [13] ResNet 56 Model 1 (44) 0.804 0.822 ResNet 56 Model 2 (45) 0.747 ResNet 56 Model 3 (46) 0.788 ResNet 56 Model 4 (47) 0.774 ResNet 56 Model 5 (48) 0.806 ResNet 56 Model 6 (49) 0.786 ResNet 56 Model 7 (50) 0.749 ResNet 56 Model 8 (51) 0.787 ResNet 110 (1.78M) [13] ResNet 110 Model 1 (44) 0.782 0.824 ResNet 110 Model 2 (45) 0.814 ResNet 110 Model 3 (46) 0.744 ResNet 110 Model 4 (47) 0.800 ResNet 110 Model 5 (48) 0.801 ResNet 110 Model 6 (49) 0.787 ResNet 110 Model 7 (50) 0.781 ResNet 110 Model 8 (51) 0.758 ResNet 152 (3.51M) [13] ResNet 152 Model 1 (44) 0.775 0.815 ResNet 152 Model 2 (45) 0.783 ResNet 152 Model 3 (46) 0.781 ResNet 152 Model 4 (47) 0.747 ResNet 152 Model 5 (48) 0.771 ResNet 152 Model 6 (49) 0.767 ResNet 152 Model 7 (50) 0.772 ResNet 152 Model 8 (51) 0.797
The results in Table 2 demonstrate the effectiveness of our genetic algorithm-based weight merging method. For ResNet 56 models, baseline test accuracies ranged from 0.747 to 0.806, improving to 0.822 after merging. This highlights our method’s ability to combine the strengths of individual models into a superior final model. Similarly, ResNet 110 models had baseline accuracies between 0.744 and 0.814, with the merged model achieving 0.824. For ResNet 152 models, baseline accuracies ranged from 0.747 to 0.783, and the merged model reached 0.815. Despite the larger complexity of ResNet 152, our method effectively enhanced model performance.
These results underscore the potential of our approach to enhance neural network performance. The consistent improvement in test accuracy across different architectures (ResNet 56, ResNet 110, and ResNet 152) demonstrates the versatility and scalability of our method. The hierarchical merging process ensures the final model benefits from the strengths of individual models. By optimizing weight combinations through selection, crossover, and mutation, we create a high-performance model that outperforms individual counterparts. This approach is particularly useful where computational resources are limited or a single robust model is preferred.

### v2

@@ 1 Introduction
Training models in parallel rather than sequentially offers efficiency and practicality. Multiple models can learn diverse patterns quickly, and merging their weights can achieve superior performance faster [14, 2, 10]. This approach is particularly advantageous in distributed learning environments, facilitating scalable and efficient training across multiple GPUs or devices [2, 22].
@@ 3.1.3 Initialization
The weights of each individual $\boldsymbol{\theta}_{i}$ are initialized as an element-wise linear combination of the weights from two pre-trained models, $\boldsymbol{\theta}_{1}$ and $\boldsymbol{\theta}_{2}$:
@@
Here, $\alpha$ is a random scalar drawn from a uniform distribution between 0 and 1. This ensures that each individual’s weights are a unique blend of the two parent models, with the combination applied element-wise to preserve the detailed characteristics of both models.
@@ 3.1.5 Selection
Next, the fitness $F(\boldsymbol{\theta})$ of each individual $\boldsymbol{\theta}$ in the tournament set $\mathcal{T}$ is evaluated using the previously defined fitness function. After evaluating the fitness of all individuals in the tournament set, the individual with the highest fitness is selected as a parent:
@@ 3.1.6 Crossover
Crossover is performed on pairs of selected parents to generate offspring. This process combines the genetic material (weights) of two parents to produce a new individual (offspring), thereby promoting genetic diversity. Given two parents $\boldsymbol{\theta}_{a}$ and $\boldsymbol{\theta}_{b}$, an offspring $\boldsymbol{\theta}_{\text{child}}$ is created by taking a weighted combination of the parents’ weights. The crossover is performed element-wise for each weight:
@@ 4.1 Experimental Settings
We utilized the CIFAR-10 dataset [18]. The CIFAR-10 training dataset was split into 45,000 training and 5,000 validation sets. The validation set was used during the model validation and merging process. We use CNN models: ResNet [12], Xception [1], and DenseNet [15] without pre-trained weights and without data augmentation. These models were trained using a batch size of 256 over 50 epochs with the Adam optimizer [17], using a learning rate of 0.01.
Each model followed the same architecture and hyperparameters to ensure consistency and a fair comparison. After training, we applied our MeGA approach to merge the weights of these models into a single set of weights. The performance of the merged model was then evaluated and compared against the individual models. For all experiments, we used the following MeGA hyperparameters: a population size of 20, 20 generations, 4 parents per generation, and a mutation rate of 0.02.
4.2 CIFAR-10 Results
The results of the image classification experiments on the CIFAR-10 dataset [18] demonstrate the effectiveness of our MeGA.
Dataset Model (# of Params.) Baseline Baseline Acc. Final Test Acc. (Random Seed) (Single Model) (Merged Model) ResNet 56 (0.86M) [12] Model 1 (56) 0.920 0.945 Model 2 (57) 0.942 SVHN [12] ResNet 110 (1.78M) [12] Model 1 (56) 0.801 0.816 (10) Model 2 (57) 0.814 ResNet 152 (3.51M) [12] Model 1 (56) 0.783 0.819 Model 2 (57) 0.781 ResNet 56 (0.86M) [12] Model 1 (56) 0.830 0.862 Model 2 (57) 0.860 Caltech-101 [12] ResNet 110 (1.78M) [12] Model 1 (56) 0.801 0.816 (10) Model 2 (57) 0.814 ResNet 152 (3.51M) [12] Model 1 (56) 0.783 0.819 Model 2 (57) 0.781 ResNet 56 (0.86M) [12] Model 1 (56) 0.830 0.862 Model 2 (57) 0.860 SVHN [12] ResNet 110 (1.78M) [12] Model 1 (56) 0.801 0.816 (10) Model 2 (57) 0.814 ResNet 152 (3.51M) [12] Model 1 (56) 0.783 0.819 Model 2 (57) 0.781
The MeGA approach led to significant improvements in test accuracies across various models as shown in the Table 2. For example, the ResNet 56 model’s accuracy improved from 0.801 to 0.822, and the ResNet 110 model reached 0.824.
Figure 2: MeGA Fitness Progression (Test Accuracy) for ResNet-56 and DenseNet 121 models over 20 generations.
The Xception model’s accuracy increased from 0.723 to 0.754, demonstrating the effectiveness of MeGA in optimizing complex architectures. Similarly, the DenseNet 121 model’s accuracy rose from 0.724 to 0.742, and the DenseNet 169 model improved from 0.733 to 0.753. Overall, the MeGA approach consistently outperformed individual models, highlighting its potential for enhancing neural network performance through genetic algorithm-based weight merging.
@@ 4.3 Extended Experiments for Multi-Models
In this section, we illustrate the application of our genetic algorithm-based weight merging method to combine multiple neural network models as depicted in Figure 3.
@@
Dataset Model (# of Params) Baseline. Baseline Test Acc. Final Test Acc. # of classes (Random Seed) (Single Model) (Merged Model) ResNet 56 (0.86M) [12] Model 1 (44) 0.804 0.822 Model 2 (45) 0.747 Model 3 (46) 0.788 Model 4 (47) 0.774 Model 5 (48) 0.806 Model 6 (49) 0.786 Model 7 (50) 0.749 Model 8 (51) 0.787 ResNet 110 (1.78M) [12] Model 1 (44) 0.782 0.824 Model 2 (45) 0.814 Model 3 (46) 0.744 CIFAR-10 [18] Model 4 (47) 0.800 (10) Model 5 (48) 0.801 Model 6 (49) 0.787 Model 7 (50) 0.781 Model 8 (51) 0.758 ResNet 152 (3.51M) [12] Model 1 (44) 0.775 0.815 Model 2 (45) 0.783 Model 3 (46) 0.781 Model 4 (47) 0.747 Model 5 (48) 0.771 Model 6 (49) 0.767 Model 7 (50) 0.772 Model 8 (51) 0.797
The hierarchical merging process starts by training eight models on the CIFAR-10 dataset [12, 18]. These models are paired and merged using a genetic algorithm, resulting in four intermediate models. These intermediate models are further merged into two higher-level models, which are finally merged into a single, robust model. This process ensures the final model combines the strengths of all eight original models. The results in Table 2 demonstrate the effectiveness of our genetic algorithm-based weight merging method. For ResNet 56 models, baseline test accuracies ranged from 0.747 to 0.806, improving to 0.822 after merging. ResNet 110 models had baseline accuracies between 0.744 and 0.814, with the merged model achieving 0.824. For ResNet 152 models, baseline accuracies ranged from 0.747 to 0.783, and the merged model reached 0.815. Despite the larger complexity of ResNet 152, our method effectively enhanced model performance.
These results highlight the potential of our approach to enhance neural network performance. The consistent improvement in test accuracy across different architectures (ResNet 56, ResNet 110, and ResNet 152) demonstrates the versatility and scalability of our method. The hierarchical merging process ensures the final model benefits from the strengths of individual models.